# Geodesics in Quantum Gravity

### v1

@@ I.2 From Quantum Gravity to Geodesic Motion
    Extremize by imposing the stationary action condition: $\delta_{x}S=0$.
@@
  • The blue path inverts the order: 3 $\rightarrow$ 1 $\rightarrow$ 2. Along this blue there is one possibility, leading to a more general equation of motion — the q–desic equation.
@@ II.1 Derivation from a Lagrangian Action
However, there is a subtlety in the final step of the procedure: Whenever we take the VEV of a non-trivial quantum operator, a projection takes place and information is lost. For example, when we take the metric VEV of equation (15), we loose information about wave functions $|\Psi\rangle$ which are not in an eigenstate of the metric operator $\hat{g}_{\mu\nu}$, or in an eigenstate of the operator $\left(\hat{g}_{\mu\alpha,\nu}+\hat{g}_{\nu\alpha,\mu}-\hat{g}_{\mu\nu,\alpha}\right)$. Clearly, we would like to reduce this loss of information to a minimum, e.g. by making one side of the equation proportional to the unit operator $\hat{1}$, since $\langle\Psi|\hat{1}|\Psi\rangle=1$, the wave function would be projected to itself and less information would be lost. Indeed, the right hand side of (15) can be made proportional to the unit operator $\hat{1}$ by invoking (6) and multiplying (15) with $\hat{g}^{\alpha\beta}$. Finally, taking the expectation value with respect to the metric degrees of freedom gives the anticipated q–desic equation
@@ II.2 Derivation from a Hamiltonian Action
If we prefer not to postpone fixing the gauge degree of freedom for the relativistic point particle, we can also work with the explicitly gauge invariant Hamiltonian action
@@
The q–desic equation (16), combined with the expectation value of the mass constraint relation (19), can be conveniently written as
@@ III.2 Selected Expectation Values
    If we could impose the normalization of the spacetime wavefunction as in quantum mechanics of a single particle, we could set $\epsilon_{0,0}$.
@@
    $\langle\hat{g}^{\prime}/\hat{g}\rangle$ $=$ $\sum_{n=1}^{\infty}\frac{\langle(2\hat{U})^{n-1}n(1-\hat{g}-r^{2}\Lambda)\rangle}{r}$ (43)
    $=$ $\frac{2G_{N}M_{0}(1+\epsilon_{1,0})}{r^{2}}-\frac{2r\Lambda}{3}+{\mathcal{O}}(\langle\hat{U}^{2}\rangle),$ (44)
@@
The VEVs presented in this subsection are parametrized by the integration constants $C_{i,j}$ which encode both the classical aspects and the quantum nature of the gravitational wave function $|\Psi\rangle$. For some of these constants, we were able to rewrite the constants, such that it is split into a classical part and a quantum part (36, 51, 52). Even though, the symbol $\epsilon_{i,j}$ suggests that the quantum part is a small correction with respect to the classical part, this does not have to be the case and for certain spacetimes the quantum part may even dominate over the classical part $\epsilon_{i,j}\geq 1$.
@@ IV.2 Energy and Angular Momentum
As consistency check, let’s recover the first and the fourth component of (57) from the conserved quantities (60 and 61). Taking the derivative of (61) with respect to the proper time and solving for $\dot{u}^{\phi}$, we get
@@ IV.4 Circular Motion of massive particles
The relations (74, 76) are important for observational purposes, since they can be translated directly to Kepler’s Third Law.
@@ VI Summary and conclusions
Looking ahead, q–desics open several avenues for further exploration. They provide a framework for probing quantum corrections to spacetime beyond the effective metric approximation, and may be extended to incorporate more general quantum states, and alternative quantum gravity models. Most importantly, q-desics give a completely new tool for observational searches for QG effects at large distances.
@@ Acknowledgements
The authors acknowledge TU Wien Bibliothek for financial support through its Open Access Funding Programme. Angel Rincon thanks Silesian University in Opava for all their support during the development of this research work.

### v2

@@ I.2 From Quantum Gravity to Geodesic Motion
    Extremize by imposing the stationary action condition after multiplying with the inverse metric: $g^{-1}\delta_{x}S=0$.
@@
  • The blue sequence changes the order: 2 $\rightarrow$ 3 $\rightarrow$ 1. It leads to a more general equation of motion – the q–desic equation.
@@ II.1 Derivation from a Lagrangian Action
However, there is a subtlety in the final step of the procedure: Whenever we take the VEV of a non-trivial quantum operator, a projection takes place and information is lost. For example, when we take the metric VEV of equation (15), we loose information about wave functions $|\Psi\rangle$ which are not in an eigenstate of the metric operator $\hat{g}_{\mu\nu}$, or in an eigenstate of the operator $\left(\hat{g}_{\mu\alpha,\nu}+\hat{g}_{\nu\alpha,\mu}-\hat{g}_{\mu\nu,\alpha}\right)$. Clearly, we would like to reduce this loss of information to a minimum, e.g. by making one side of the equation proportional to the unit operator $\hat{1}$. Since $\langle\Psi|\hat{1}|\Psi\rangle=1$, the wave function would be projected to itself and less information would be lost. Indeed, the right hand side of (15) can be made proportional to the unit operator $\hat{1}$ by invoking (6) and multiplying (15) with $\hat{g}^{\alpha\beta}$. Finally, taking the expectation value with respect to the metric degrees of freedom gives the anticipated q–desic equation
@@ II.2 Derivation from a Hamiltonian Action
If we prefer to postpone fixing the gauge degree of freedom for the relativistic point particle, we can also work with the explicitly gauge invariant Hamiltonian action
@@
The q–desic equation (16), combined with the expectation value of the mass constraint relation (19), while the latter can be conveniently written as
@@ III.2 Selected Expectation Values
    If we could impose the normalization of the spacetime wavefunction as in quantum mechanics of a single particle, we could set $\epsilon_{0,0}$ to zero.
@@
The VEVs presented in this subsection are parametrized by the integration constants $C_{i,j}$ which encode both the classical aspects and the quantum nature of the gravitational wave function $|\Psi\rangle$. For some of these constants, we were able to rewrite the constants, such that it is split into a classical part and a quantum part (36, 49, 50). Even though, the symbol $\epsilon_{i,j}$ suggests that the quantum part is a small correction with respect to the classical part, this does not have to be the case and for certain spacetimes the quantum part may even dominate over the classical part $\epsilon_{i,j}\geq 1$.
@@ IV.2 Energy and Angular Momentum
As consistency check, let’s recover the first and the fourth component of (55) from the conserved quantities (58 and 59). Taking the derivative of (59) with respect to the proper time and solving for $\dot{u}^{\phi}$, we get
@@ IV.4 Circular Motion of Massive Particles
The relations (72, 74) are important for observational purposes, since they can be translated directly to Kepler’s Third Law.
@@ VI Summary and conclusions
Looking ahead, q–desics open several avenues for further exploration. They provide a framework for probing quantum corrections to spacetime beyond the effective metric approximation, and may be extended to incorporate more general quantum states, and alternative quantum gravity models. Most importantly, q–desics give a completely new tool for observational searches for QG effects at large distances.
Note added: After the article’s first appearance on the arXiv, we became aware of three related approaches:
  1. 1.
    The impact of virtual gravitons and backreaction was studied in Dalvit and Mazzitelli (1997, 1999). The calculations use perturbation theory, expanding the metric around a semiclassical background $\tilde{g}_{\mu\nu}=g_{\mu\nu}+\kappa h_{\mu\nu}$, followed by path–integral methods to derive an effective action for both the background and a test particle. Although the techniques and language differ from our approach, it is natural to assume that this effective test–particle action corresponds to the VEV of our operator–valued action (18), i.e. $\langle\hat{S}_{H}\rangle$. Extremizing this expectation value of the average action yields the VEV of equation (23), leads to
    $\frac{d^{2}x^{\mu}}{d\lambda^{2}}\;+\;\Gamma^{\mu}_{\alpha\beta}\bigl(\langle\tilde{g}_{\gamma\delta}\rangle\bigr)\,\frac{dx^{\alpha}}{d\lambda}\,\frac{dx^{\beta}}{d\lambda}\;+\frac{1}{m}\langle\tilde{g}^{\mu\nu}\rangle\frac{\delta\Delta S_{m}}{\delta x^{\nu}}=\;0\,$,
    where $\Delta S_{m}$ contains the $\sim h_{\mu\nu}^{2}$ contributions to the point particle action. Even if $\Delta S_{m}=0$, this is not the classical geodesic equation, since the effective metric $\tilde{g}_{\mu\nu}$ contains contributions from a nonclassical spacetime state. Nevertheless, it contains less information than the q–desic equation (16), due to the subtlety explained between equations (15) and (16), where the VEV is taken before acting with the inverse metric operator. The difference with respect to the q–desic is exactly the information about the spacetime state, which can not be obtained in terms of universal loop calculations and can not be expressed in terms of the universal constants ($G,\,\hbar,\,c,\,M$).
  2. 2.
    Perturbative quantum fluctuations of the metric field have also been considered in the definition of geodesic motion with respect to the orthonormal frame of an observers worldline Parikh et al. (2021); Chawla and Parikh (2023); Bak et al. (2023, 2024); Cho and Hu (2022); Hsiang et al. (2024). It was shown that a path integral over such fluctuations produces a characteristic but $\Psi$-state dependent stochastic force $N_{\Psi}$, modifying the geodesic equations. For example, even in the Newtonian limit
    $\ddot{\vec{x}}=-\vec{\nabla}\phi+N_{\Psi}\vec{x}+\dots\quad.$ (89)
    Such an effect could in principle be identified by observing the correlations of numerous different geodesics. In contrast, the q–desic (16), can already be manifest in a single trajectory.
  3. 3.
    The definition of a spatial distance operator in the context of QG provides the starting point for another series of papers Piazza (2025); Nitti et al. (2024); Piazza and Tolley (2023); Piazza (2023). The authors do not attempt to compute a semiclassical metric; instead they focus on the spacetime state $\Psi$. They show that, under certain assumptions, the state is not peaked around the classical value and distances become nonadditive. In our approach, this nonadditivity arises as a special case of the covariance in equation (V.2). In this sense, those works anticipated several of our findings based on this non-trivial covariance.
Interestingly, in approaches 1 and 2 the deviations from classical geodesics are expected only at extremely short distances, whereas in approach 3 the dominant deviations occur at very large distances. The q–desics derived here allow for both short– and long–range modifications; see Fig. 2.
@@ Acknowledgements
The authors acknowledge TU Wien Bibliothek for financial support through its Open Access Funding Programme. Ángel Rincón thanks Silesian University in Opava for all their support during the development of this research work.